# Accretion of Generalized Chaplygin Gas onto Cosmologically Coupled Black Holes

### v1

@@ I Introduction
While the cosmological constant remains the simplest and most widely adopted model for dark energy, a variety of dynamical models have been proposed to address the theoretical and observational challenges. Among these, the Generalized Chaplygin Gas (GCG) has attracted attention due to its ability to interpolate between dust-like matter at early times and a dark energy–dominated regime at late times [kamenshchik2001alternative, Bilic:2001cg, bento2002generalized].
The GCG model not only provides a compelling observational framework to account for the universe’s accelerated expansion and to address the dark matter problem [bento2003wmap, aurich2018concordance, barreiro2008wmap], but also offers promising avenues for addressing fundamental issues such as the cosmological constant problem [bertolami2023seeding], the Hubble tension [yang2019dawn], the Swampland conjectures [bertolami2024sitter], and inflation through a model inspired on the GCG [Bertolami:2006zg].
@@
Some studies have already explored dark energy and the GCG in the context of compact objects, such as in Refs. [Mazur:2001fv, bertolami2005chaplygin, gorini2009more]. Other works have also investigated dark energy and the GCG in accretion scenarios, such as in Refs. [babichev2005accretion, rodrigues2012accretion], although there remains substantial room for further development in this area.
Classical accretion theory has been instrumental in describing the infall of baryonic matter and the associated electromagnetic signatures [bondi1952spherically, petterson1980variations, michel1972accretion, porth2017black]. However, in the context of a universe dominated by a dark energy, it becomes imperative to consider whether and how black holes may accrete this non-standard component.
The accretion of dark energy, particularly when modeled as a perfect fluid with negative pressure, introduces novel theoretical possibilities. In scenarios involving phantom energy, that satisfies the condition $p+\rho<-1$, it has been shown that the accretion process could lead to a decrease in black hole mass, a counterintuitive result that challenges conventional views of black hole growth and thermodynamics [babichev2004black]. However, despite the consistency of such approaches, certain aspects are often neglected, largely due to the considerable complexity involved in developing a comprehensive accretion model.
@@
For this purpose, we adopt the McVittie metric [mcvittie1933mass]. This spacetime solution has been the subject of ongoing debate for nearly eighty years regarding its physical consistency in describing black holes embedded in an expanding universe. However, recent studies have demonstrated that the McVittie metric not only provides a consistent framework for modeling black holes that are asymptotically FLRW [kaloper2010mcvittie, lake2011more, nandra2012effect, nandra2012effect1, faraoni2015cosmological], but also serves as an important foundational model for the development of more sophisticated and realistic metrics to this end [gao2011black].
@@
This paper is organized as follows: In Section II, We analyze the McVittie metric, with an emphasis on the effective mass associated with this geometry and on the time-dependent solutions of the apparent horizons. In Section III, we use a perturbative analysis that has already been developed that accounts for backreaction effects to study accretion within the McVittie spacetime. In Section IV, we apply the results developed in Section III to the GCG, focusing on two distinct analyses: one corresponding to a matter-dominated universe and the other to a de Sitter universe. Finally, in Section V, we present our conclusions and discuss the implications of the results.
@@
II The McVittie metric and the GCG
The McVittie metric can be expressed in several equivalent forms; however, in this work, we adopt the formulation presented in Ref. [kaloper2010mcvittie]. The line element is given by
@@ II The McVittie metric and the GCG
$-3H_{0}^{2}(t)$ $=$ $8\pi\left(T_{0}^{\text{ }0}\right)^{\left(0\right)},$ (2a)
$-3H_{0}^{2}-\frac{2\dot{H}_{0}}{\sqrt{1-\frac{2M_{0}}{r}}}$ $=$ $8\pi\left(T_{1}^{\text{ }1}\right)^{\left(0\right)}.$ (2b)
If we define $\left(T_{\nu}^{\text{ }\mu}\right)^{\left(0\right)}$ such that $\left(T_{0}^{\text{ }0}\right)^{\left(0\right)}=-\rho^{\left(0\right)}$ and $\left(T_{1}^{\text{ }1}\right)^{\left(0\right)}=p^{\left(0\right)}$, where $\rho^{\left(0\right)}$ and $p^{\left(0\right)}$ are the energy density and isotropic pressure, respectively, Eq. (2a) can be readily recognized as the Friedmann equation. This indicates that the presence of a black hole does not affect the homogeneity of the energy density distribution. However, Eq. (2b) reveals that the same is not true for the pressure, which acquires a spatial dependence due to the black hole’s influence.
@@
where $f\left(t\right)$ is an arbitrary function of time. The authors in Refs. [nandra2012effect, nandra2012effect1] have shown that for a matter-dominated universe, it is appropriate to take $f\left(t\right)=\rho^{\left(0\right)}$, such that the pressure can be written as
@@
where $A>0$ and $0<\alpha\leq 1$. The Chaplygin gas corresponds to the choice $\alpha=1$ [kamenshchik2001alternative, bento2002generalized, bento2003wmap]. In the framework of FLRW cosmology, the relativistic energy conservation associated to the equation of state (6) leads to an evolving energy density:
@@
Other key aspects of the McVittie metric relevant to our analysis are the effective mass and the apparent horizon. The effective mass of the black hole can be computed using the Misner-Sharp mass function $M_{MS}^{\left(0\right)}$ [misner1964relativistic], defined by
@@
which, using Eqs. (1) and (2a), results in
@@
The apparent horizon of the black hole, denoted $r_{h}^{\left(0\right)}$, can be determined from the condition $g_{00}^{\left(0\right)}=0$, which, using Eq. (1), leads to the following cubic equation:
$H_{0}^{2}\left(r_{h}^{\left(0\right)}\right)^{3}-r_{h}^{\left(0\right)}+2M_{0}=0.$ (11)
As shown in Refs. [nickalls1993new, faraoni2012making] the solutions of the cubic equation above can be expressed as:
$r_{1}\left(t\right)$ $=$ $\frac{2}{\sqrt{3}H_{0}\left(t\right)}\sin\left(\frac{\arcsin\left(3\sqrt{3}M_{0}H_{0}\left(t\right)\right)}{3}\right),$ (12a)
@@
As discussed in Ref. [faraoni2012making], the solution $r_{3}$ can be discarded as it is unphysical, since it is always negative. The solution $r_{1}$ can be interpreted as the apparent horizon of the black hole, as it reduces to the Schwarzschild radius $r_{1}\rightarrow 2M_{0}$ in the limit $H_{0}\rightarrow 0$. Similarly, $r_{2}$ can be interpreted as the cosmological horizon, as it approaches $r_{2}\rightarrow 1/H_{0}$ when $M_{0}\rightarrow 0$. Henceforth we will denote $r_{1}$ by $r_{b}$ and $r_{2}$ by $r_{c}$.
Furthermore, the solutions given by Eqs. (12a) and (12) impose the following conditions: for $3\sqrt{3}M_{0}H_{0}<1$, the horiozons $r_{b}$ and $r_{c}$ are distinct; when $3\sqrt{3}M_{0}H_{0}=1$, the black hole horizon coincides with the cosmological horizon; and when $3\sqrt{3}M_{0}H_{0}>1$, no horizons exist and the black hole becomes a naked singularity.
@@
III Perturbatitve approach to backreaction
The process of accretion within the framework of the McVittie metric has been investigated in previous works [gao2011black, gao2008does, guariento2012realistic]. The approach adopted in these studies primarily consists in promoting the mass parameter $M_{0}$ in Eq. (1) to a time-dependent function.
In the present work, we assume that the accretion of matter induces backreaction effects on the geometry. To investigate this backreaction, we employ a perturbative method as proposed in Ref. [babichev2012backreaction].
We begin by assuming that the backreaction induces both a temporal and radial dependence of the mass, as well as an inhomogeneity in the Hubble parameter. Consequently, the metric (1) becomes:
$ds^{2}=-\left(1-\frac{2M\left(t,r\right)}{r}-H\left(t,r\right)^{2}r^{2}\right)dt^{2}-\frac{2rH\left(t,r\right)}{\sqrt{1-\frac{2M\left(t,r\right)}{r}}}dtdr+\frac{dr^{2}}{1-\frac{2M\left(t,r\right)}{r}}+r^{2}d\Omega_{2}^{2}.$ (14)
@@ III Perturbatitve approach to backreaction
$8\pi\mathcal{T}_{0}^{\text{ }0}$ $=$ $-3H^{2}-2rHH^{\prime}-\frac{2M^{\prime}}{r^{2}},$ (15a)
$8\pi\mathcal{T}_{0}^{\text{ }1}$ $=$ $2rH\dot{H}+\frac{2\dot{M}}{r^{2}},$ (15b)
$8\pi\mathcal{T}_{1}^{\text{ }0}$ $=$ $-\frac{2\dot{M}}{\left(r-2M\right)^{2}},$ (15c)
$8\pi\mathcal{T}_{1}^{\text{ }1}$ $=$ $-\frac{1}{r^{2}\left(r-2M\right)^{2}}\left[3r^{2}H^{2}\left(r-2M\right)^{2}\right$.
@@
One also finds that the Einstein tensors $G_{2}^{\text{ }2}=G_{3}^{\text{ }3}\neq 0$; however, these components yield no independent equations, as they follow from Eqs. (15a)-(15c).
@@
where $g_{\mu\nu}$ is the metric associated with the line element (14), then Eqs. (15a) and (15b) can be rewritten as
@@
[…]
@@
$\mathcal{M}_{MS}\left(t,r\right)$ $=$ $M_{i}+\int_{t_{0ac}}^{t}\mathcal{A}\left(\bar{t},r_{0ac}\right)~d\bar{t}+\int_{r_{0ac}}^{r_{b}}\mathcal{B}\left(t,\bar{r}\right)~d\bar{r},$
@@
A solution similar to Eq. (LABEL:M_solution) appears in Refs. [babichev2012backreaction, campos2025dynamical]. However, in the present context, we are dealing with a slightly more general scenario, since both the accretion term and the energy density are allowed to depend on time. Moreover, it is worth highlighting that working with the McVittie metric allows us to obtain solutions analogous to that of Refs. [babichev2012backreaction, campos2025dynamical], but expressed in a physical time coordinate rather than in an Eddington–Finkelstein–type time coordinate.
The expression for $\mathcal{M}_{MS}$ reveals the contribution of the accretion of matter, described by $\mathcal{T}_{\nu}^{\text{ }\mu}$, to the initial effective mass $M_{i}$. The first integral in (LABEL:M_solution) corresponds to the contribution from the evolution of the accreting matter flux, evaluated in the region bounded by $r_{0ac}$ over the interval $t-t_{0ac}$. The second integral represents the time-dependent contribution from the energy density contained between the initial apparent horizon before accretion, $r_{0ac}$ , and the apparent horizon displaced due to accretion, $r_{b}$.
With the solution (LABEL:M_solution) at hand, we can now effectively compute the effects of accretion. We only need to evaluate the integrals in Eq.(LABEL:M_solution) in terms of the energy–momentum tensors. We shall assume that the tensors $\mathcal{T}_{\nu}^{\ \mu}$ satisfying Einstein’s equations (15) correspond to a perfect fluid, given by
@@
However, an evident difficulty immediately arises: the metric components are expressed in terms of the unknown functions $M\left(t,r\right)$ and $H\left(t,r\right)$, which appear in Eq. (14). To address this, we shall employ the perturbative approach used in Ref. [babichev2012backreaction]. We assume that the effects of the accreting fluid are much smaller than those of the background fluid, so that the energy–momentum tensor of the accreting fluid can be evaluated in the background metric $g_{\mu\nu}^{\left(0\right)}$ given by (1). Consequently, the tensor $\mathcal{T}_{\nu}^{\text{ }\mu}$ and the solution (LABEL:M_solution) will be considered as a first-order perturbation, namely
@@
$\left(T_{0}^{\text{ }0}\right)^{\left(1\right)}\left(t,r_{b}\right)$ $=$ $-\frac{3H_{0}^{2}\left(t\right)}{8\pi}+\frac{\left(\rho_{ac}+p_{ac}\right)H_{0}^{2}\left(t\right)\left(r_{b}^{\left(0\right)}\left(t\right)\right)^{2}}{v-2H_{0}^{2}\left(t\right)\left(r_{b}^{\left(0\right)}\left(t\right)\right)^{2}}$ (28)
$+p_{ac}$
@@
Under this perturbative approximation, Eq. (LABEL:M_solution) expressed in terms of the Eqs. (28) and (29) is exact at the initial time $t=t_{0ac}$, such that
@@
We may then analyze the integrals appearing in Eq. (LABEL:M_solution) using the definitions from Eqs. (19) and (20). We are going to define the first integral in Eq. (LABEL:M_solution), after substituting Eq. (29), as
@@
Since our goal is to study the accretion of the GCG, before evaluating the second integral in Eq. (LABEL:M_solution), we anticipate that $\rho_{ac}$ and $p_{ac}$ with which we will work depend only on time. Evidently, the same assumption holds for the radial velocity of the fluid, $v$, once our goal is to evaluate the accretion at a specific point, the black hole horizon, i.e., $v(t,r)=v(t,r^{(1)}_{b})$. Consequently, $\left(T_{0}^{\text{ }0}\right)^{\left(1\right)}$ given by Eq. (28), will depend solely on time, and therefore Eq. (20) becomes
$\int_{r_{0ac}}^{r_{b}^{\left(1\right)}}\mathcal{B}\left(t,\bar{r}\right)~d\bar{r}$ $=$ $-4\pi\left(T_{0}^{\text{ }0}\left(t\right)\right)^{\left(1\right)}\int_{r_{0ac}}^{r_{b}^{\left(1\right)}}\bar{r}^{2}~d\bar{r}=$ (32)
$=$ $\frac{H_{1}^{2}\left(t\right)}{2}\left(r_{b}^{\left(1\right)}\right)^{3}+G\left(t\right),$
@@
Finally, the solution given in Eq. (LABEL:M_solution) can be expressed as
@@
This equation is structurally analogous to the cubic Eq. (11). Its physically meaningful solutions, in agreement with (12a) and (12), are given by
@@
$r_{c}^{\left(1\right)}\left(t\right)$ $=$ $\frac{1}{H_{1}\left(t\right)}\cos\left(\frac{\arcsin\left(3\sqrt{3}M_{1}\left(t\right)H_{1}\left(t\right)\right)}{3}\right)$
$-\frac{1}{\sqrt{3}H_{1}\left(t\right)}\sin\left(\frac{\arcsin\left(3\sqrt{3}M_{1}\left(t\right)H_{1}\left(t\right)\right)}{3}\right)$
. (39b)
@@
Moreover, the conditions required for the solutions in Eqs. (39) to be real are similar to those discussed in the previous section for Eqs. (12a) and (12), now for $M_{0}\rightarrow M_{1}\left(t\right)$ e $H_{0}\left(t\right)\rightarrow H_{1}\left(t\right)$.
@@
IV Applications
IV.1 Dark energy accretion

### v2

@@ 1 Introduction
While the cosmological constant remains the simplest and most widely adopted model for dark energy, a variety of dynamical models have been proposed to address the theoretical and observational challenges. Among these, the Generalized Chaplygin Gas (GCG) has attracted attention due to its ability to interpolate between dust-like matter at early times and a dark energy-dominated regime at late times [8, 9, 10].
The GCG model not only provides a compelling observational framework to account for the universe’s accelerated expansion and to address the dark matter problem [11, 12, 13], but also offers promising avenues for addressing fundamental issues such as the cosmological constant problem [14], the Hubble tension [15], the Swampland conjectures [16], and inflation through a model inspired by the GCG [17].
@@
Some studies have already explored dark energy and the GCG in the context of compact objects, such as in Refs. [18, 19, 20]. Other works have also investigated dark energy and the GCG in accretion scenarios, such as in Refs. [21, 22], although there remains substantial room for further development in this area.
Classical accretion theory has been instrumental in describing the infall of baryonic matter and the associated electromagnetic signatures [23, 24, 25, 26]. However, in the context of a universe dominated by a dark energy, it becomes imperative to consider whether and how black holes may accrete this non-standard component.
The accretion of dark energy, particularly when modeled as a perfect fluid with negative pressure, introduces novel theoretical possibilities. In scenarios involving phantom energy, that satisfies the condition $p+\rho<-1$, it has been shown that the accretion process could lead to a decrease in black hole mass, a counterintuitive result that challenges conventional views of black hole growth and thermodynamics [27]. However, despite the consistency of such approaches, certain aspects are often neglected, largely due to the considerable complexity involved in developing a comprehensive accretion model.
@@
Backreaction effects have been explored in steady accretion models where the self-gravity of the accreting fluid modifies the background geometry; for example, the fully relativistic treatment in [28] examines how the gravitational influence of matter affects accretion compared to the Bondi limit, and Ref. [29] studies steady critical accretion effects on the characteristics of the sonic points.
@@
For this purpose, we adopt the McVittie metric [30]. This spacetime solution has been the subject of ongoing debate for nearly eighty years regarding its physical consistency in describing black holes embedded in an expanding universe. However, recent studies have demonstrated that the McVittie metric not only provides a consistent framework for modeling black holes that are asymptotically FLRW [31, 32, 33, 34, 35], but also serves as an important foundational model for the development of more sophisticated and realistic metrics to this end [36].
@@
This paper is organized as follows: In Section 2 we analyze the McVittie metric, with an emphasis on the effective mass associated with this geometry and on the time-dependent solutions of the apparent horizons. In Section 3, we use a perturbative analysis that has already been developed that accounts for backreaction effects to study accretion within the McVittie spacetime. In Section 4, we apply the results developed in Section 3 to the GCG, focusing on two distinct analyses: one corresponding to a matter-dominated universe and the other to a de Sitter universe. Finally, in Section 5, we present our conclusions and discuss the implications of the results.
@@
2 The McVittie metric and the GCG
The McVittie metric can be expressed in several equivalent forms; however, in this work, we adopt the formulation presented in Ref. [31]. The line element is given by
@@ 2 The McVittie metric and the GCG
$8\pi\left(T_{0}^{\text{ }0}\right)^{\left(0\right)}$ $=-3H_{0}^{2}(t),$ (2)
$8\pi\left(T_{1}^{\text{ }1}\right)^{\left(0\right)}$ $=-3H_{0}^{2}-\frac{2\dot{H}_{0}}{\sqrt{1-\frac{2M_{0}}{r}}}.$ (3)
If we define $\left(T_{\nu}^{\text{ }\mu}\right)^{\left(0\right)}$ such that $\left(T_{0}^{\text{ }0}\right)^{\left(0\right)}=-\rho^{\left(0\right)}$ and $\left(T_{1}^{\text{ }1}\right)^{\left(0\right)}=p^{\left(0\right)}$, where $\rho^{\left(0\right)}$ and $p^{\left(0\right)}$ are the energy density and isotropic pressure, respectively, Eq. (2) can be readily recognized as the Friedmann equation. This indicates that the presence of a black hole does not affect the homogeneity of the energy density distribution. However, Eq. (3) reveals that the same is not true for the pressure, which acquires a spatial dependence due to the black hole’s influence.
@@
where $f\left(t\right)$ is an arbitrary function of time. The authors in Refs. [33, 34] have shown that for a matter-dominated universe, it is appropriate to take $f\left(t\right)=\rho^{\left(0\right)}$, such that the pressure can be written as
@@
where $A>0$ and $0<\alpha\leq 1$. The Chaplygin gas corresponds to the choice $\alpha=1$ [8, 10, 11]. In the framework of FLRW cosmology, the relativistic energy conservation associated to the equation of state (7) leads to an evolving energy density:
@@
Other key aspects of the McVittie metric relevant to our analysis are the effective mass and the apparent horizon. The effective mass of the black hole can be computed using the Misner-Sharp mass function $M_{MS}^{\left(0\right)}$ [37], defined by
@@
which, using Eqs. (1) and (2), results in
@@
Since the McVittie spacetime is dynamical according to Eq. (11), a global event horizon cannot, in general, be identified without knowledge of the full future development of the spacetime. It is therefore more appropriate to characterize black hole boundaries quasi-locally in terms of apparent horizons.
In a given spacelike hypersurface, an apparent horizon is defined as the outermost marginally outer trapped surface (MOTS). One may equivalently characterize the apparent horizon in terms of the optical scalars—namely, the expansion $\Theta$, the twist $\omega$, and the shear $\sigma$. In this framework, the apparent horizon is specified by the conditions $\Theta_{+}=0$, $\Theta_{-}<0$, $\omega=0$, and $\sigma=0$, where $\Theta_{+}=0$ corresponds to the vanishing expansion of the outgoing null congruence, while $\Theta_{-}<0$ ensures that the ingoing null congruence remains converging [28].
For a spherically symmetric spacetime given by the metric (1), the symmetry two-spheres have area $\mathcal{S}=4\pi r^{2}$ and, in the coordinates used here, the outgoing and ingoing expansions take the form
$\Theta_{\pm}=\frac{2}{r}\frac{dr}{ds_{\pm}}\propto\sqrt{1-\frac{2M_{0}}{r}}\left(H_{0}r\pm\sqrt{1-\frac{2M_{0}}{r}}\right).$ (12)
Therefore, the marginally trapped condition $\Theta_{+}=0$ yields
$H_{0}^{2}\left(r_{h}^{\left(0\right)}\right)^{3}-r_{h}^{\left(0\right)}+2M_{0}=0,$ (13)
where $r_{h}^{\left(0\right)}$ represents the geometrical locus of the apparent horizon.
As shown in Refs. [38, 39] the solutions of the cubic equation above can be expressed as:
$r_{1}\left(t\right)$ $=\frac{2}{\sqrt{3}H_{0}\left(t\right)}\sin\left(\frac{\arcsin\left(3\sqrt{3}M_{0}H_{0}\left(t\right)\right)}{3}\right),$ (14)
@@
As discussed in Ref. [39], the solution $r_{3}$ can be discarded as it is unphysical, since it is always negative. The solution $r_{1}$ can be interpreted as the apparent horizon of the black hole, as it reduces to the Schwarzschild radius $r_{1}\rightarrow 2M_{0}$ in the limit $H_{0}\rightarrow 0$. Similarly, $r_{2}$ can be interpreted as the cosmological horizon, as it approaches $r_{2}\rightarrow 1/H_{0}$ when $M_{0}\rightarrow 0$. Henceforth we will denote $r_{1}$ by $r_{b}$ and $r_{2}$ by $r_{c}$.
Furthermore, the solutions given by Eqs. (14)-(2) impose the following conditions: for $3\sqrt{3}M_{0}H_{0}<1$, the horizons $r_{b}$ and $r_{c}$ are distinct; when $3\sqrt{3}M_{0}H_{0}=1$, the black hole horizon coincides with the cosmological horizon; and when $3\sqrt{3}M_{0}H_{0}>1$, no horizons exist and the black hole becomes a naked singularity.
@@
3 Perturbative approach to backreaction
The process of accretion within the framework of the McVittie metric has been investigated in previous works [36, 40, 41]. The approach adopted in these studies primarily consists in promoting the mass parameter $M_{0}$ in Eq. (1) to a time-dependent function.
In the present work, we assume that the accretion of matter induces backreaction effects on the geometry. To investigate this backreaction, we employ a perturbative method as proposed in Ref. [42]. We do it so by assuming that accretion induces both a temporal and radial dependence of the mass, as well as an inhomogeneity in the Hubble parameter. Consequently, the metric (1) becomes:
$ds^{2}$ $=$ $-\left(1-\frac{2M\left(t,r\right)}{r}-H\left(t,r\right)^{2}r^{2}\right)dt^{2}-\frac{2rH\left(t,r\right)}{\sqrt{1-\frac{2M\left(t,r\right)}{r}}}dtdr$ (18)
$+$ $\frac{dr^{2}}{1-\frac{2M\left(t,r\right)}{r}}+r^{2}d\Omega_{2}^{2}.$
@@ 3 Perturbative approach to backreaction
$8\pi\mathcal{T}_{0}^{\text{ }0}$ $=-3H^{2}-2rHH^{\prime}-\frac{2M^{\prime}}{r^{2}},$ (19)
$8\pi\mathcal{T}_{0}^{\text{ }1}$ $=2rH\dot{H}+\frac{2\dot{M}}{r^{2}},$ (20)
$8\pi\mathcal{T}_{1}^{\text{ }0}$ $=-\frac{2\dot{M}}{\left(r-2M\right)^{2}},$ (21)
$8\pi\mathcal{T}_{1}^{\text{ }1}$ $=-\frac{1}{r^{2}\left(r-2M\right)^{2}}\left[3r^{2}H^{2}\left(r-2M\right)^{2}\right$. (22)
@@
One also finds that the Einstein tensors $G_{2}^{\text{ }2}=G_{3}^{\text{ }3}\neq 0$; however, these components yield no independent equations, as they follow from Eqs. (19)-(21).
@@
where $g_{\mu\nu}$ is the metric associated with the line element (18), then Eqs. (19) and (20) can be rewritten as
@@
$\dot{\mathcal{M}}_{MS}\left(t,r\right)=\mathcal{A}\left(t,r\right)$ (25)
@@
$\mathcal{M}_{MS}\left(t,r\right)=M_{i}+\int_{t_{0ac}}^{t}\mathcal{A}\left(\bar{t},r_{0ac}\right)~d\bar{t}+\int_{r_{0ac}}^{r_{b}}\mathcal{B}\left(t,\bar{r}\right)~d\bar{r},$ (28)
@@
A solution similar to Eq. (28) appears in Refs. [42, 43]. However, in the present context, we are dealing with a slightly more general scenario, since both the accretion term and the energy density are allowed to depend on time. Moreover, it is worth highlighting that working with the McVittie metric allows us to obtain solutions analogous to that of Refs. [42, 43], but expressed in a physical time coordinate rather than in an Eddington–Finkelstein–type time coordinate.
The expression for $\mathcal{M}_{MS}$ reveals the contribution of the accretion of matter, described by $\mathcal{T}_{\nu}^{\text{ }\mu}$, to the initial effective mass $M_{i}$. The first integral in (28) corresponds to the contribution from the evolution of the accreting matter flux, evaluated in the region bounded by $r_{0ac}$ over the interval $t-t_{0ac}$. The second integral represents the time-dependent contribution from the energy density contained between the initial apparent horizon before accretion, $r_{0ac}$ , and the apparent horizon displaced due to accretion, $r_{b}$.
With the solution (28) at hand, we can now effectively compute the effects of accretion. We only need to evaluate the integrals in Eq. (28) in terms of the energy–momentum tensors. We shall assume that the tensors $\mathcal{T}_{\nu}^{\ \mu}$ satisfying Einstein’s equations (19)-(22) correspond to a perfect fluid, given by
@@
However, an evident difficulty immediately arises: the metric components are expressed in terms of the unknown functions $M\left(t,r\right)$ and $H\left(t,r\right)$, which appear in Eq. (18). To address this, we shall employ the perturbative approach used in Ref. [42]. We assume that the effects of the accreting fluid are much smaller than those of the background fluid, so that the energy–momentum tensor of the accreting fluid can be evaluated in the background metric $g_{\mu\nu}^{\left(0\right)}$ given by (1). Consequently, the tensor $\mathcal{T}_{\nu}^{\text{ }\mu}$ and the solution (28) will be considered as a first-order perturbation, namely
@@
$\left(T_{0}^{\text{ }0}\right)^{\left(1\right)}\left(t,r_{b}\right)=-\frac{3H_{0}^{2}\left(t\right)}{8\pi}+\frac{\left(\rho_{ac}+p_{ac}\right)H_{0}^{2}\left(t\right)\left(r_{b}^{\left(0\right)}\left(t\right)\right)^{2}}{v-2H_{0}^{2}\left(t\right)\left(r_{b}^{\left(0\right)}\left(t\right)\right)^{2}}+p_{ac}$ (35)
@@
Under this perturbative approximation, Eq. (28) expressed in terms of the Eqs. (35) and (36) is exact at the initial time $t=t_{0ac}$, such that
@@
We may then analyze the integrals appearing in Eq. (28) using the definitions from Eqs. (26) and (27). We are going to define the first integral in Eq. (28), after substituting Eq. (36), as
@@
Since our goal is to study the accretion of the GCG, before evaluating the second integral in Eq. (28), we anticipate that $\rho_{ac}$ and $p_{ac}$ with which we will work depend only on time. Evidently, the same assumption holds for the radial velocity of the fluid, $v$, once our goal is to evaluate the accretion at a specific point, the black hole horizon, i.e., $v(t,r)=v(t,r^{(1)}_{b})$. Consequently, $\left(T_{0}^{\text{ }0}\right)^{\left(1\right)}$ given by Eq. (35), will depend solely on time, and therefore Eq. (27) becomes
$\int_{r_{0ac}}^{r_{b}^{\left(1\right)}}\mathcal{B}\left(t,\bar{r}\right)~d\bar{r}=-4\pi\left(T_{0}^{\text{ }0}\left(t\right)\right)^{\left(1\right)}\int_{r_{0ac}}^{r_{b}^{\left(1\right)}}\bar{r}^{2}~d\bar{r}=\frac{H_{1}^{2}\left(t\right)}{2}\left(r_{b}^{\left(1\right)}\right)^{3}+G\left(t\right),$ (39)
@@
Finally, the solution given in Eq. (28) can be expressed as
@@
This equation is structurally analogous to the cubic Eq. (13). Its physically meaningful solutions, in agreement with (14) and (2), are given by
@@
$r_{c}^{\left(1\right)}\left(t\right)$ $=\frac{1}{H_{1}\left(t\right)}\cos\left(\frac{\arcsin\left(3\sqrt{3}M_{1}\left(t\right)H_{1}\left(t\right)\right)}{3}\right)$ (48)
$-\frac{1}{\sqrt{3}H_{1}\left(t\right)}\sin\left(\frac{\arcsin\left(3\sqrt{3}M_{1}\left(t\right)H_{1}\left(t\right)\right)}{3}\right).$
@@
Moreover, the conditions required for the solutions in Eqs. (LABEL:r1b)-(48) to be real are similar to those discussed in the previous section for Eqs. (14) and (2), now for $M_{0}\rightarrow M_{1}\left(t\right)$ and $H_{0}\left(t\right)\rightarrow H_{1}\left(t\right)$.
@@
4 Applications
4.1 Dark energy accretion